OpenStudy (anonymous):

With a principal investment of $19,200, which account will have the greatest value after 5 years? simple interest: I = P • r • t interest compounded annually is A = P (1 + r)t interest compounded quarterly: A = P (1 + r/4 )4t A. 3.6% with interest compounded annually B. 3.8% in a simple interest account C. 3.4% with interest compounded annually D. 3.2% with interest compounded quarterly

OpenStudy (raffle_snaffle):

What do you think?

OpenStudy (raffle_snaffle):

Describe to me what the difference is between simple and compound interest?

Directrix (directrix):

Take the first option and use it to crank out the interest with the given data of P = 19 200 and t = 5 years Option A Test: A. 3.6% with interest compounded annually interest compounded annually is A = P (1 + r)^t A = 19 200 ( 1 + .036)^5 A = $ 4535 interest Total amount after 5 years is 19 200 + 4 535 = ? After doing that, move on to check the total amounts for other 3 options. @OpenStudier0